# Hook length biases in ordinary and $t$-regular partitions

### v1

@@ Abstract.
In this article, we study hook lengths of ordinary partitions and $t$-regular partitions. We establish hook length biases for the ordinary partitions and motivated by them we find a few interesting hook length biases in $2$-regular partitions. For a positive integer $k$, let $p_{(k)}(n)$ denote the number of hooks of length $k$ in all the partitions of $n$. We prove that $p_{(k)}(n)\geq p_{(k+1)}(n)$ for all $n\geq 0$ and $n\neq k+1$; and $p_{(k)}(k+1)-p_{(k+1)}(k+1)=-1$ for $k\geq 2$. For integers $t\geq 2$ and $k\geq 1$, let $b_{t,k}(n)$ denote the number of hooks of length $k$ in all the $t$-regular partitions of $n$. We find generating functions of $b_{t,k}(n)$ for certain values of $t$ and $k$. Exploring hook length biases for $b_{t,k}(n)$, we observe that in certain cases biases are opposite to the biases for ordinary partitions. We prove that $b_{2,2}(n)\geq b_{2,1}(n)$ for all $n>4$, whereas $b_{2,2}(n)\geq b_{2,3}(n)$ for all $n\geq 0$. Our proofs are both combinatorial and analytic in nature. We also propose some conjectures on biases among $b_{t,k}(n)$.
@@ 1. Introduction and statement of results
A Young diagram of a partition $(\lambda_{1},\lambda_{2},\ldots,\lambda_{r})$ is a left-justified array of boxes with $i$-th row (from the top) having $\lambda_{i}$ boxes. For example, the Young diagram of the partition $(5,4,2,2,1)$ is shown in Figure 1 (left). A box in the $i$-th row and $j$-th column of the Young diagram of a partition $\lambda$ can be denoted by $(i,j)$. The hook length of a box $(i,j)$ is the number of the boxes directly right and directly below the box, plus 1 for the box itself. For example, see Figure 1 (right) for the hook lengths of each box in the Young diagram of the partition $(5,4,2,2,1)$.
@@
Hook lengths of partitions have important connection with representation theory of the symmetric groups $S_{n}$ and $GL_{n}(\mathbb{C})$. Hook lengths also appear in the Seiberg-Witten theory of random partitions, which gives Nekrasov-Okounkov formula for arbitrary powers of Euler’s infinite product in terms of hook numbers. For more in these directions, see e.g. [5, 9, 10, 11, 13]. Other than the ordinary partition function, hook lengths have been studied for several restricted partition functions, for example, partitions into odd parts, partitions into distinct parts, partitions into odd and distinct parts, self conjugate partitions and doubled distinct partitions, see e.g. [1, 3, 4, 7, 8, 12]. In this article, we study the total number of hooks of fixed length in $t$-regular partitions along with the number of hooks of fixed length in ordinary partitions. We establish hook length biases in ordinary partitions and motivated by them we find a few hook length biases for $t$-regular partitions.
@@
Table 1 exhibits the values of $p_{(k)}(n)$ for some values of $k$ and $n$. Since the Young diagram of a partition $\lambda$ of $n\leq k-1$ can have at most $k-1$ boxes, there is no hook of length $k$ for $\lambda$. Hence, $p_{(k)}(n)=0$ for all $n$ satisfying $11\leq n\leq k-$ 1. Also, observe that the values in every column of Table 1 are in non-increasing fashion, except from the last non-zero value in the column. That is, counting in all the partitions of $n$, the number of hooks of length $k$ is at least the number of hooks of length $k+1$, with one exception of $n=k+1$. We confirm this in the following theorem.
@@
In this article, we obtain the generating functions of $b_{t,k}(n)$ for certain values of $t$ and $k$. First, we find the generating function of $b_{t,1}(n)$ for all $t\geq 2$.
@@
We propose Conjecture 1.6 on the basis of data obtained by enumerating partitions, since it is difficult to explicitly derive the generating functions of $b_{2,k}(n)$ for higher values of $k$.
@@
[…]
@@
[…]
@@ 2. Hook length biases in ordinary partitions
In this section we prove Theorem 1.1. We first discuss some prerequisites. We define non-negativity of a series $S(q):=\sum_{n\geq 0}s_{n}q^{n}$, if $s_{n}\geq 0$, for all $n$. We use the notation $S(q)\succcurlyeq 0$ to denote the non-negativity.
@@
Clearly, $\Psi_{n}$ is a one-one map. Thus, $|\mathcal{P}^{1}(n)|\leq|\mathcal{P}^{1}(n+1)|$ and therefore $a(n)\leq a(n+1)$. ∎
@@
$=\mathcal{K}_{k}(q)+\frac{\mathcal{L}_{k}(q)}{(1-q^{k^{2}+k})(q;q)_{\infty}},$ (2.4)
@@ 3. Generating functions for $b_{t,k}(n)$
In this section, we derive the generating functions of $b_{t,1}(n)$ for all $t\geq 2$. We also derive the generating functions for $b_{3,2}(n)$ and $b_{4,2}(n)$. First, we introduce some notations that would be useful in this section. For any part $s$ of a partition $\lambda$, we denote by $m_{\lambda}(s)$ the multiplicity of $s$ in $\lambda$, i.e., the number of times $s$ appears as a part in $\lambda$. For a positive integer $j$, let $\ell_{>j}(\lambda)$ denote the number of parts of $\lambda$ which are greater than $j$. By $\ell_{1}(\lambda)$ (respectively $\ell_{2}(\lambda)$) we denote the number of parts $\lambda_{i}$ of $\lambda$ with $\lambda_{i}-\lambda_{i+1}=1$ (respectively $\lambda_{i}-\lambda_{i+1}=2$), where we take $\lambda_{i}=0$ if $i>\ell(\lambda)$. Let $\delta_{P}$ denote the Kronecker delta symbol, which is equal to 1 if property $P$ is true and 0 otherwise.
@@
$\frac{\partial}{\partial z}\bigg{|}_{z=1}\mathcal{B}^{1}_{3,2}(z;q)$ $=\prod_{\begin{subarray}{c}n\geq 1\\ 3\nmid n\end{subarray}}\frac{1}{1-q^{n}}\left(q^{2}+\sum_{\begin{subarray}{c}n\geq 2\\ 3\nmid n\end{subarray}}\left(q^{n}+q^{2n}\right)\right)$ (3.9)
@@
$\frac{\partial}{\partial z}\bigg{|}_{z=1}\mathcal{B}^{2}_{4,2}(z;q)=$ $\prod_{\begin{subarray}{c}n\geq 1\\ 4\nmid n\end{subarray}}\frac{1}{1-q^{n}}\left(\prod_{n\geq 0}(1-q^{4n+1}q^{4n+3})\right)\times$ (3.22)
@@ 4. Hook length biases in 2-regular partitions
In this section, we give proofs of Theorems 1.4 and 1.5.
@@
For $n>4$, let $\mathcal{B}_{2}(n)$ denote the set of all 2-regular or odd partitions of $n$. Also, let $\mathcal{R}_{n}\subset\mathcal{B}_{2}(n)$ be the set of partitions having at least one part of multiplicity more than 1; $\mathcal{S}_{n}\subset\mathcal{B}_{2}(n)$ be the set of distinct partitions with smallest part 1; and $\mathcal{T}_{n}\subset\mathcal{B}_{2}(n)$ be the set of distinct partitions with smallest part greater than 1. Then $\mathcal{B}_{2}(n)=\mathcal{R}_{n}\cup\mathcal{S}_{n}\cup\mathcal{T}_{n}$.
Let $\lambda\in\mathcal{R}_{n}$. If a part $\lambda_{i}$ of $\lambda$ occurs at least twice then the rows corresponding to $\lambda_{i}$ in the Young diagram of $\lambda$ will have two hooks of length 2 and one hook of length 1. For every other part of $\lambda$ which is greater than one, there is one hook of length 1 and at least one hook of length 2. For part of size 1 (if present as a part in $\lambda$), there is a hook of length 1 and there may or may not be a hook of length 2. Considering all the cases, we conclude that the number of hooks of length 1 in $\lambda$ is not more than the number of hooks of length 2 in $\lambda$.
Define a map $\Phi_{n}:\mathcal{S}_{n}\rightarrow\mathcal{R}_{n}$ by
@@
$\Phi_{n}\left((\lambda_{1}=n-1,1)\right):=\left\{\begin{array}[]{lll}((\frac{n-3}{3})^{3},1^{3})&\text{if}\ n\equiv 0\pmod{3};\\ ((\frac{n-4}{3})^{3},1^{4})&\text{if}\ n\equiv 1\pmod{3};\\ ((\frac{n-2}{3})^{3},1^{2})&\text{if}\ n\equiv 2\pmod{3}.\end{array}\right.$
Clearly, $\Phi_{n}$ is a one-one map. For any $\lambda\in\mathcal{S}_{n}$, the number of hooks of length 1 in $\lambda$ is exactly one more than the number of hooks of length 2 in $\lambda$. For any $\lambda\in\Phi_{n}(\mathcal{S}_{n})$, the number of hooks of length 1 in $\lambda$ is exactly one less than the number of hooks of length 2 in $\lambda$. Thus, every $\lambda\in\mathcal{S}_{n}$, which is having one extra hook of length 1 than the hooks of length 2, corresponds to $\Phi_{n}(\lambda)\in\mathcal{R}_{n}$, which is having one extra hook of length 2 than the hooks of length 1. Hence, the number of hooks of length 1 in all the partitions in $\mathcal{R}_{n}\cup\mathcal{S}_{n}$ is less than the number of hooks of length 2 in all the partitions on $\mathcal{R}_{n}\cup\mathcal{S}_{n}$.
If $\lambda\in\mathcal{T}_{n}$, then every part of $\lambda$ corresponds to exactly one hook of length 1 and one hook of length 2 in the Young diagram of $\lambda$. Therefore, in this case the number of hooks of length 1 in $\lambda$ is same as the number of hooks of length 2 in $\lambda$.
The above all cases conclude that in all the 2-regular partitions of $n$, the number of hooks of length 1 is less than the number of hooks of length 2. ∎
@@
5. Values of $b_{t,k}(n)$ for $1\leq n\leq k$
In this section, we prove Theorems 1.7 and 1.8. First, we recall another form of representation of a partition given by
$\lambda=(\lambda^{m_{1}}_{1},\lambda^{m_{2}}_{2},\ldots,\lambda^{m_{r}}_{r}),$
@@ 6. Concluding Remarks
The motivating questions arising from this study can be greatly generalized. We see in Theorem 1.1 that the number of hooks of length $k$ is at least the number of hooks of length $k+1$ in all partitions of $n$, for all $n$ except $n=k+1$. But in Theorem 1.4, we find that this bias reverses for $k=1$ in the case of 2-regular partitions. In the direction of Conjecture 6.1 and following Theorem 1.4, it would be interesting to study the relationship between $b_{t,1}(n)$ and $b_{t,2}(n)$, for $t\geq 3$. Also, we see in Theorem 1.5 that the bias for 2-regular partitions remains the same as in the ordinary partitions for $k=2$. Apart from Conjecture 1.6, it would also be interesting to study the relationship between $b_{t,k}(n)$ and $b_{t,k+1}(n)$, for $t\geq 3$ and $k\geq 2$.

### v2

@@ Abstract.
In this article, we study hook lengths of ordinary partitions and $t$-regular partitions. We establish hook length biases for the ordinary partitions and motivated by them we find a few interesting hook length biases in $2$-regular partitions. For a positive integer $k$, let $p_{(k)}(n)$ denote the number of hooks of length $k$ in all the partitions of $n$. We prove that $p_{(k)}(n)\geq p_{(k+1)}(n)$ for all $n\geq 0$ and $n\neq k+1$; and $p_{(k)}(k+1)-p_{(k+1)}(k+1)=-1$ for $k\geq 2$. For integers $t\geq 2$ and $k\geq 1$, let $b_{t,k}(n)$ denote the number of hooks of length $k$ in all the $t$-regular partitions of $n$. We find generating functions of $b_{t,k}(n)$ for certain values of $t$ and $k$. Exploring hook length biases for $b_{t,k}(n)$, we observe that in certain cases biases are opposite to the biases for ordinary partitions. We prove that $b_{2,2}(n)\geq b_{2,1}(n)$ for all $n>4$, whereas $b_{2,2}(n)\geq b_{2,3}(n)$ for all $n\geq 0$. We also propose some conjectures on biases among $b_{t,k}(n)$.
@@ 1. Introduction and statement of results
A Young diagram of a partition $(\lambda_{1},\lambda_{2},\ldots,\lambda_{r})$ is a left-justified array of boxes with $i$-th row (from the top) having $\lambda_{i}$ boxes. For example, the Young diagram of the partition $(5,4,2,2,1)$ is shown in Figure 1 (left). The hook length of a box in a Young diagram is the sum of the number of the boxes directly right to it, the number of boxes directly below it and 1 (for the box itself). For example, see Figure 1 (right) for the hook lengths of each box in the Young diagram of the partition $(5,4,2,2,1)$.
@@
Hook lengths of partitions have important connection with representation theory of the symmetric groups $S_{n}$ and $GL_{n}(\mathbb{C})$. Hook lengths also appear in the Seiberg-Witten theory of random partitions, which gives Nekrasov-Okounkov formula for arbitrary powers of Euler’s infinite product in terms of hook numbers. For more in these directions, see e.g. [5, 9, 10, 11, 13]. Other than the ordinary partition function, hook lengths have also been studied for several restricted partition functions, for example, partitions into odd parts, partitions into distinct parts, partitions into odd and distinct parts, self conjugate partitions and doubled distinct partitions, see e.g. [1, 3, 4, 7, 8, 12]. In this article, we study the total number of hooks of fixed length in $t$-regular partitions along with the number of hooks of fixed length in ordinary partitions. We establish hook length biases in ordinary partitions and motivated by them we find a few hook length biases for $t$-regular partitions.
@@
Table 1 exhibits the values of $p_{(k)}(n)$ for some values of $k$ and $n$. Since the Young diagram of a partition $\lambda$ of $n\leq k-1$ can have at most $k-1$ boxes, there is no hook of length $k$ for $\lambda$. Hence, $p_{(k)}(n)=0$ for all $n$ satisfying $11\leq n\leq k-$ 1. Also, observe that the values in every column of Table 1 are in non-increasing order, except from the last non-zero value in the column. That is, counting in all the partitions of $n$, the number of hooks of length $k$ is at least the number of hooks of length $k+1$, with one exception of $n=k+1$. We confirm this in the following theorem.
@@
Very recently, Craig et al. [4] obtained the generating functions for $b_{2,k}(n)$, for any $k$. In this article, using the ideas of Ballantine et al. [3], we obtain the generating functions of $b_{t,k}(n)$ for certain new values of $t$ and $k$. First, we find the generating function of $b_{t,1}(n)$ for all $t\geq 2$.
@@
$b_{2,k}(k)$ =$=\left\{\begin{array}[]{lll}\frac{k+1}{2}&\text{if}\ k\equiv 1\pmod{2};\\ \frac{k}{2}&\text{if}\ k\equiv 0\pmod{2},\end{array}\right.$ (1.11)
$b_{2,k}(k+1)$ =$=\left\{\begin{array}[]{lll}1&\text{if}\ k\equiv 1\pmod{2};\\ 2&\text{if}\ k\equiv 0\pmod{2}.\end{array}\right.$ (1.14)
@@
    $b_{t,k}(k+1)$ =$=\left\{\begin{array}[]{lll}\frac{(t-1)k+t}{t}&\text{if}\ k\equiv 0\pmod{t};\vspace{.12cm}\\ \frac{(t-1)k+r}{t}&\text{if}\ k\equiv r\pmod{t}\ \text{for}\ 1\leq r\leq t-2;\vspace{.12cm}\\ \frac{(t-1)k-1}{t}&\text{if}\ k\equiv t-1\pmod{t}.\end{array}\right.$ (1.22)
@@
The rest of this paper is organized as follows. In Section 2 we prove Theorem 1.1 which gives certain hook length biases in ordinary partition functions. In Section 3 we derive the generating functions of $b_{t,1}(n)$ for all $t\geq 2$, namely we prove Theorem 1.2. We also prove Theorem 1.3 which gives the generating functions for $b_{3,2}(n)$ and $b_{4,2}(n)$. In Section 4 we prove the hook length biases in $2$-regular partitions, namely Theorems 1.4 and 1.5. In Section 5 we prove Theorems 1.7 and 1.8 which give values of $b_{t,k}(k)$ and $b_{t,k}(k+1)$. In Section 6 we make some concluding remarks and state some problems for future works.
@@ 2. Hook length biases in ordinary partitions
In this section we prove Theorem 1.1. We first discuss some prerequisites. We say that a series $S(q):=\sum_{n\geq 0}s_{n}q^{n}$ is non-negative if $s_{n}\geq 0$, for all $n$. We use the notation $S(q)\succcurlyeq 0$ to denote the non-negativity.
@@
Clearly, $\Psi_{n}$ is a one-to-one map. Thus, $|\mathcal{P}^{1}(n)|\leq|\mathcal{P}^{1}(n+1)|$ and therefore $a(n)\leq a(n+1)$. ∎
@@
Note that (2.1) is also an immediate consequence of Lemma 2.1.
@@
$=\mathcal{K}_{k}(q)+\frac{\mathcal{L}_{k}(q)}{(1-q^{k^{2}+k})(q^{2};q)_{\infty}},$ (2.4)
@@ 3. Generating functions for $b_{t,k}(n)$
In this section, we derive the generating functions of $b_{t,1}(n)$ for all $t\geq 2$. We also derive the generating functions for $b_{3,2}(n)$ and $b_{4,2}(n)$. First, we introduce some notations that would be useful in this section. For any part $s$ of a partition $\lambda$, we denote by $m_{\lambda}(s)$ the multiplicity of $s$ in $\lambda$, i.e., the number of times $s$ appears as a part in $\lambda$. For a positive integer $j$, let $\ell_{>j}(\lambda)$ denote the number of part sizes greater than $j$ in the partition $\lambda$. By $\ell_{1}(\lambda)$ (respectively $\ell_{2}(\lambda)$) we denote the number of parts $\lambda_{i}$ of $\lambda$ with $\lambda_{i}-\lambda_{i+1}=1$ (respectively $\lambda_{i}-\lambda_{i+1}=2$), where we take $\lambda_{i}=0$ if $i>\ell(\lambda)$. Let $\delta_{P}$ denote the Kronecker delta symbol, which is equal to 1 if property $P$ is true and 0 otherwise.
@@
$\frac{\partial}{\partial z}\bigg{|}_{z=1}\mathcal{B}^{1}_{3,2}(z;q)$ $=\prod_{\begin{subarray}{c}n\geq 1\\ 3\nmid n\end{subarray}}\frac{1}{1-q^{n}}\left(q^{2}+\sum_{\begin{subarray}{c}n\geq 2\\ 3\nmid n\end{subarray}}\left(q^{n}+q^{2n}\right)\right)$
@@
$\frac{\partial}{\partial z}\bigg{|}_{z=1}\mathcal{B}^{2}_{4,2}(z;q)=$ $\prod_{\begin{subarray}{c}n\geq 1\\ 4\nmid n\end{subarray}}\frac{1}{1-q^{n}}\left(\prod_{n\geq 0}(1-q^{4n+1}q^{4n+3})\right)\times$
@@ 4. Hook length biases in 2-regular partitions
In this section, we give proofs of Theorems 1.4 and 1.5. First, we recall another form of representation of a partition given by
$\lambda=(\lambda^{m_{1}}_{1},\lambda^{m_{2}}_{2},\ldots,\lambda^{m_{r}}_{r}),$
@@
For $n>4$, let $\mathcal{B}_{2}(n)$ denote the set of all 2-regular or odd partitions of $n$. Also, let $\mathcal{R}_{n}\subset\mathcal{B}_{2}(n)$ be the set of partitions having at least one part of multiplicity more than 1; $\mathcal{S}_{n}\subset\mathcal{B}_{2}(n)$ be the set of distinct partitions with smallest part 1; and $\mathcal{T}_{n}\subset\mathcal{B}_{2}(n)$ be the set of distinct partitions with smallest part greater than 1. Then $\mathcal{B}_{2}(n)=\mathcal{R}_{n}\cup\mathcal{S}_{n}\cup\mathcal{T}_{n}$. Given a partition $\lambda\in\mathcal{B}_{2}(n)$, let $h_{k}(\lambda)$ denote the number of hooks of length $k$ in the partition $\lambda$.
Case 1: Let $\lambda\in\mathcal{R}_{n}$. If a part $\lambda_{i}$ of $\lambda$ occurs at least twice then the rows corresponding to $\lambda_{i}$ in the Young diagram of $\lambda$ will have two hooks of length 2 and one hook of length 1. For every other part of $\lambda$ which is greater than one, there is one hook of length 1 and at least one hook of length 2. For part of size 1 (if present as a part in $\lambda$), there is a hook of length 1 and there may or may not be a hook of length 2. Considering all the cases, we conclude that the number of hooks of length 1 in $\lambda$ is not more than the number of hooks of length 2 in $\lambda$. Hence, $h_{2}(\lambda)\geq h_{1}(\lambda)$ for all $\lambda\in\mathcal{R}_{n}$.
Case 2: Let $\lambda\in\mathcal{S}_{n}$. Clearly, the number of hooks of length 1 in $\lambda$ is exactly one more than the number of hooks of length 2 in $\lambda$. That is, $h_{1}(\lambda)=1+h_{2}(\lambda)$ for all $\lambda\in\mathcal{S}_{n}$.
Case 3: Let $\lambda\in\mathcal{T}_{n}$. Then, every part of $\lambda$ corresponds to exactly one hook of length 1 and one hook of length 2 in the Young diagram of $\lambda$. Therefore, in this case the number of hooks of length 1 in $\lambda$ is same as the number of hooks of length 2 in $\lambda$. Thus, $h_{1}(\lambda)=h_{2}(\lambda)$ for all $\lambda\in\mathcal{T}_{n}$.
In order to establish the required hook lengths bias, we next show that the extra hook of length 1 coming from Case 2 will be adjusted in Case 1. For this, we define a map $\Phi_{n}:\mathcal{S}_{n}\rightarrow\mathcal{R}_{n}$ by
@@
$\Phi_{n}\left((\lambda_{1}=n-1,1)\right):=\left\{\begin{array}[]{lll}((\frac{n-2}{2})^{2},1^{2})&\text{if}\ n\equiv 0\pmod{4};\vspace{.12cm}\\ ((\frac{n}{2})^{2})&\text{if}\ n\equiv 2\pmod{4}.\end{array}\right.$
Clearly, $\Phi_{n}$ is a one-to-one map. For any $\lambda\in\Phi_{n}(\mathcal{S}_{n})$, the number of hooks of length 1 in $\lambda$ is exactly one less than the number of hooks of length 2 in $\lambda$. Thus, $h_{1}(\lambda)=h_{2}(\lambda)-1$ for all $\lambda\in\Phi_{n}(\mathcal{S}_{n})$.
Combining all the cases and the fact that $\Phi_{n}$ is one-to-one, we have
$b_{2,2}(n)$ $=\sum_{\lambda\in\mathcal{B}_{2}(n)}h_{2}(\lambda)$
$=\sum_{\lambda\in\mathcal{R}_{n}}h_{2}(\lambda)+\sum_{\lambda\in\mathcal{S}_{n}}h_{2}(\lambda)+\sum_{\lambda\in\mathcal{T}_{n}}h_{2}(\lambda)$
$=\sum_{\lambda\in\mathcal{R}_{n}\setminus\Phi_{n}(\mathcal{S}_{n})}h_{2}(\lambda)+\sum_{\lambda\in\Phi_{n}(\mathcal{S}_{n})}h_{2}(\lambda)+\sum_{\lambda\in\mathcal{S}_{n}}h_{2}(\lambda)+\sum_{\lambda\in\mathcal{T}_{n}}h_{2}(\lambda)$
$\geq\sum_{\lambda\in\mathcal{R}_{n}\setminus\Phi_{n}(\mathcal{S}_{n})}h_{1}(\lambda)+\sum_{\lambda\in\Phi_{n}(\mathcal{S}_{n})}(h_{1}(\lambda)+1)+\sum_{\lambda\in\mathcal{S}_{n}}(h_{1}(\lambda)-1)+\sum_{\lambda\in\mathcal{T}_{n}}h_{1}(\lambda)$
$=\sum_{\lambda\in\mathcal{B}_{2}(n)}h_{1}(\lambda)$
$=b_{2,1}(n).$
This completes the proof of the theorem. ∎
@@
5. Values of $b_{t,k}(k)$ and $b_{t,k}(k+1)$
In this section, we prove Theorems 1.7 and 1.8.
@@ 6. Concluding Remarks
This study gives rise to many interesting questions. We see in Theorem 1.1 that the number of hooks of length $k$ is at least the number of hooks of length $k+1$ in all partitions of $n$, for all $n$ except $n=k+1$. But in Theorem 1.4, we find that this bias reverses for $k=1$ in the case of 2-regular partitions. In the direction of Conjecture 6.1 and following Theorem 1.4, it would be interesting to study the relationship between $b_{t,1}(n)$ and $b_{t,2}(n)$, for $t\geq 3$. Also, we see in Theorem 1.5 that the bias for 2-regular partitions remains the same as in the ordinary partitions for $k=2$. Apart from Conjecture 1.6, it would also be interesting to study the relationship between $b_{t,k}(n)$ and $b_{t,k+1}(n)$, for $t\geq 3$ and $k\geq 2$.
@@
7. Acknowledgement
The authors thank the referee for many valuable comments. The second author gratefully acknowledge the Department of Science and Technology, government of India, for the Core Research Grant (CRG/2021/00314) of SERB.